OpenStudy (anonymous):

what is an easy way to convert fractions to decimals??? including mixed fractions

OpenStudy (anonymous):

divide... by long division but best by calculator.

OpenStudy (asnaseer):

You might find this site useful: http://www.mathsisfun.com/converting-fractions-decimals.html

OpenStudy (campbell_st):

numerator divided by the denominator or know some basics 1/2 = 0.5 1/4 = 0.25 1/5 = 0.2 1/10 = 0.1

OpenStudy (anonymous):

thanks everyone!

OpenStudy (asnaseer):

yw